# Read Between the Layers: Leveraging Intra-Layer Representations for Rehearsal-Free Continual Learning with Pre-Trained Models

### v1

@@ Abstract
We address the Continual Learning (CL) problem, where a model has to learn a sequence of tasks from non-stationary distributions while preserving prior knowledge as it encounters new experiences. With the advancement of foundation models, CL research has shifted focus from the initial learning-from-scratch paradigm to the use of generic features from large-scale pre-training. However, existing approaches to CL with pre-trained models only focus on separating the class-specific features from the final representation layer and neglect the power of intermediate representations that capture low- and mid-level features naturally more invariant to domain shifts. In this work, we propose LayUP, a new class-prototype-based approach to continual learning that leverages second-order feature statistics from multiple intermediate layers of a pre-trained network. Our method is conceptually simple, does not require any replay buffer, and works out of the box with any foundation model. LayUP improves over the state-of-the-art on four of the seven class-incremental learning settings at a considerably reduced memory and computational footprint compared with the next best baseline. Our results demonstrate that fully exhausting the representational capacities of pre-trained models in CL goes far beyond their final embeddings. The code will be made publicly available upon acceptance.
@@ 1 Introduction
Figure 1: Overview of LayUP: Enriching last layer representations of pre-trained models with features from intermediate layers leads to better calibrated similarity measures for prototype-based classification and more robustness to domain gaps in the continual learning setting.
@@
Continual Learning (CL) is a subfield of machine learning that focuses on enabling models to learn from a stream of data and adapt to novel concepts without forgetting previously acquired knowledge [45, 6]. While traditional works on CL primarily focus on the learning-from-scratch paradigm, the introduction of large foundation models has initiated a growing interest in developing CL methods upon the powerful representations resulting from large-scale pre-training. Continual learning with pre-trained models builds on the assumption that a large all-purpose feature extractor provides strong knowledge transfer capabilities and great robustness to catastrophic forgetting [39] during incremental adaptation to downstream tasks.
@@
Such methods for downstream adaptation can be categorized into two general strategies [61, 40]: (i) Carefully fine-tuning the pre-trained representations (full body adaptation), or (ii) keeping the pre-trained backbone fixed while learning additional parameters (such as prompts [31, 35, 22], adapters [50, 19, 20, 5, 7], or other methods [25, 38, 34]). Both strategies have in common that they do not rely on rehearsal from past data to perform well, unlike many of the top-performing CL strategies used for models trained from scratch. Despite their merits, most approaches to CL with pre-trained models that fall under one of those strategies update a fully connected classification head (linear probe) via iterative gradient methods during training and inference, which is subject to task-recency bias [37, 54, 62] and catastrophic forgetting [69, 48]. Class-prototype methods [72, 21, 40] have been established as a viable alternative, as they average extracted features directly, which is not only computationally efficient but also superior to CL methods applied to the trainable linear probe.
@@
Despite their effectiveness, contemporary class-prototype methods for CL use only the features obtained by the last layer of the backbone for class-prototype construction. However, as the discrepancy between the pre-training and fine-tuning domain increases, the high-level last-layer features might not generalize well enough to provide a good class separability in the target domain anymore. The challenge is therefore to find a prototype-based classifier that leverages the representations of a pre-trained feature extractor holistically, ideally at comparably low memory and computational cost.
@@
In this work, we hypothesize that representations are most effectively leveraged from multiple layers to construct a classifier based on first-order (class-prototypes) and second-order (Gram matrix) feature statistics. Such a strategy is inspired by works on Neural Style Transfer [13, 23], which involve applying the artistic style of one image to the content of another image, a problem that can be viewed through the lens of domain adaptation [32]. To achieve neural style transfer, a model has to disentangle image information into content and style (e.g., textures, patterns, colors). While content information is expressed in terms of activations of features at different layers (which translates to calculating intra-layer class-prototypes in our case), style information is expressed as correlations between features of different layers (which corresponds to us calculating intra-layer Gram matrices).
@@
We thus propose Intra-Layer Universal Prototypes (LayUP), a class-prototype method for CL that is based on the aforementioned strategy to disentangle style and content information of images at multiple layers of a pre-trained network. An overview of our method is given in Fig. 1. LayUP obtains rich and fine-grained information about images to perform a more informed classification that is less sensitive to domain shifts. We additionally experiment with different parameter-efficient fine-tuning strategies to further refine the intra-layer representations obtained by LayUP. Across several CL benchmarks and learning settings, our method provides improvements and reduces the gap to upper bound performance by up to 67% (Stanford Cars-196 in the CIL setting) compared with the next best baseline while reducing its memory and compute requirements by 81% and up to 90%, respectively.
@@
Our contributions are threefold: (1) We present and examine in detail the CL strategy with pre-trained models and show why it benefits from using intra-layer representations for classification. We further demonstrate the advantages of leveraging cross-correlations between intra-layer and inter-layer features to decorrelate class-prototypes. (2) Building on our insights, we propose LayUP, a novel class-prototype approach to CL that leverages second-order statistics of features from multiple layers of a pre-trained model. Inspired by prior works [71, 40], we experiment with different methods for parameter-efficient model tuning and extend our method towards first session adaptation. Our final approach is conceptually simple, light in memory and compute, and works out-of-the-box with any transformer model. (3) We report performance improvements with a pre-trained Vit-B/16 [12] backbone on several benchmarks both in the Class-Incremental Learning (CIL) and in the more challenging Online Continual Learning (OCL) setting and show that LayUP is especially effective under large distributional shifts and in the low-data regime. Our results highlight the importance of taking a deeper look at the intermediate layers of pre-trained models to better leverage their powerful representations, thus identifying promising directions for CL in the era of large foundation models.
@@ 2.1 Continual learning
Traditional approaches to CL consider training a model from scratch on a sequence of tasks, while striking a balance between adapting to the currently seen task and maintaining high performance on previous tasks. They can be broadly categorized into regularization [27, 67, 33, 1, 11], which selectively restrict parameter updates, (pseudo-)rehearsal [51, 4, 2, 47], which recover data either from a memory buffer or from a generative model, and dynamic architectures [53, 66, 55], which allocate disjoint parameter spaces for each task.
@@ 2.2 Continual learning with pre-trained models
Leveraging the powerful representations from large foundation models for downstream continual learning has shown to not only facilitate knowledge transfer, but also to increase robustness against forgetting [49, 43]. L2P [65] and DualPrompt [64] outperform traditional CL baselines by training a pool of learnable parameters (i.e., prompts) that are used as queries to guide a Vision Transformer (ViT) [12] towards adapting to downstream tasks. They serve as the basis for several more recent prompt learning methods such as S-Prompt [63], DAP [24], and CODA-Prompt [56], which further increase performance.
Contrary to prompt learning methods that keep the backbone parameters fixed while updating a small set of additional parameters, some parallel works explore methods for making careful adjustments to the backbone parameters. SLCA [69] uses a small learning rate for updates to the ViT backbone, while training a linear probing layer with a larger learning rate. L2 [57] applies regularization to the self-attention parameters of a ViT during continual fine-tuning. First Session Adaptation (FSA) [44] makes adjustments to the parameters of the backbone only during first task training.
As an alternative to the methods above that adjust the feature space via cross-entropy loss and backpropagation, constructing prototypes directly from the pre-trained features can avoid catastrophic forgetting during CL without storing data. Prior works [21, 72] find that a conceptually simple Nearest Mean Classifier (NMC) outperforms various prompt learning methods. ADAM [72] combines NMC with FSA and concatenates the features from the initial pre-trained ViT and the adapted parameters. Other methods leverage second-order feature statistics, i.e., covariance information, for class-prototype accumulation: FSA+LDA [44] applies an incremental version of linear discriminant analysis to a pre-trained ResNet encoder [14]. RanPAC [40] uses high-dimensional random feature projections to decorrelate class-prototypes of a pre-trained ViT backbone.
@@ 3 Preliminaries
3.1 Continual learning problem setting
We consider a feature extractor $\phi(\cdot)$ composed of $L$ consecutive layers and a sequence of training datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$, where the $t$th task $\mathcal{D}_{t}=\{({\bm{x}}_{t,n},y_{t,n})\}_{n=1}^{N_{t}}$ contains pairs of input samples ${\bm{x}}_{t,n}\in\mathcal{X}_{t}$ and their ground-truth labels $y_{t,n}\in\mathcal{Y}_{t}$. Let $\mathcal{Y}_{t}$ denote the label space of the $t$th task and $\mathcal{Y}=\bigcup_{t=1}^{T}\mathcal{Y}_{t}$. The total number of classes seen in $\mathcal{D}$ (i.e., the number of unique elements in $\mathcal{Y}$) is denoted with $C$. We represent the $d_{l}$-dimensional encoded features from the $l$th layer of model $\phi$ from some test image ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$ with unknown label during inference as $\phi_{l}({\bm{x}})\in{\mathbb{R}}^{d_{l}}$. We consider rehearsal-free CL, where historical data can not be fetched for replay. Experiments in Sec. 5 are conducted under Class-Incremental Learning (CIL) and Online Continual Learning (OCL) settings which prohibit the access of the model to task identifiers during testing (notably, our method does not need any task-specific information during training, either). OCL is a more tightened definition of CIL towards each datum in the stream of tasks being observed only once.
@@ 3.2 Class-prototype methods for CL with pre-trained models
To leverage the powerful representations from large-scale pre-training while overcoming the limitations of full body adaptation and linear probing, class-prototype methods construct representatives for each class from the last layer features of a pre-trained model. The most straightforward class-prototype method is the Nearest Mean Classifier (NMC), which aggregates per-class training sample features via averaging (denoted as $\overline{{\bm{c}}}_{y}$ for each class $y$) and, during inference, selects for each test sample’s feature vector the class with the smallest Euclidean distance [21] or highest cosine similarity [72] with its class-prototype. Considering the cosine similarity measure and some ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, the predicted class label is obtained by:
@@
Although Eq. 1 and Eq. 2 are defined with respect to the maximum number of classes $C$ after observing all $T$ tasks, they can be applied after seeing any task $t\leq T$ or any sample $n\leq N_{t}$. When denoting the extracted features of some input datum ${\bm{x}}_{t,n}\in\mathcal{D}_{t}$ as ${\bm{f}}_{t,n}=\phi_{L}({\bm{x}}_{t,n})$ and considering ${\bm{F}}\in\mathbb{R}^{N\times d_{L}}$ as concatenated row-vector features ${\bm{f}}_{t,n}$ of all $N$ training samples, Eq. 3 will be reduced to ${\bm{G}}={\bm{F}}^{T}{\bm{F}}$, which corresponds to the original definition of a Gram matrix as used in the closed-form ridge estimator [18].
4 LayUP: Intra-Layer Universal Prototypes
4.1 Motivation
@@
denote the concatenated input features of some input sample ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, extracted from the last $k$ layers of the pre-trained model $\phi(\cdot)$. Such features can be, e.g., class embeddings ([CLS]) of a transformer [59] encoder or flattened feature maps of a ResNet [14] encoder. This yields a modified estimator
@@
We first motivate our approach by showing that intra-layer representations capture expressive class statistics for prototype-based classification. For the split CIFAR-100 ($C=100$) and ImageNet-R ($C=200$) datasets, we construct one prototype-based classifier using Eq. 2 per layer $l\in\{1,\dots,L\}$ and count the total number of classes that layer $l$ predicts better than any other layer $l^{\prime}\neq l$. Results are provided in Fig. 1(a). We observe that especially the last four intermediate layers (8–11) account for a top performance of between 25% and 54% of all classes, indicating that the representations obtained by intermediate layers can generally add meaningful knowledge to better separate classes at different hierarchical levels.
(a) Intra-layer performance
(b) LayUP (shared) vs. averaging over layer-wise similarity scores (separate)
Figure 2: (a) Comparison of classification performance of layers $(1,\dots,L)$ for two different pre-training paradigms of a ViT-B/16 [12] backbone and split ImageNet-R and CIFAR-100 datasets. Each bar indicates the absolute count of classes for which a classifier constructed using Eq. 2 for the respective layer yields highest classification accuracy compared to all other layers.
@@
(b) Average accuracy (%) over CL training on split ImageNet-R and CIFAR-100 datasets following phase B of Alg. 1: Comparison of LayUP implementations for different values of $k$ using Eq. 5 versus averaging over separate similarity scores for each layer using Eq. 2.
@@
There are two intuitive ways to integrate intra-layer representations into the class-prototype classifier: Averaging over $k$ separate classifiers per Eq. 2 or concatenating representations from the last $k$ layers to obtain shared class prototypes per Eq. 5. Fig. 1(b) shows average accuracies for two split datasets and different values of $k$. For each $k$, using shared class-specific information as per Eq. 5 is superior to averaging over separate classifiers, showing that the correlations across layers, which are captured in the shared Gram matrix (cf. Eq. 6), add meaningful information to enhance class separability and thus improve performance. Although this leads us to choose higher memory demands compared with a last layer classifier and averaging over layer-wise classifiers, we will show in Sec. 4.3 that our method is still considerably lighter in memory and computation than other competitive class-prototype methods for CL.
@@
Interestingly, the performance of the last layer classifier is roughly the same compared with averaging over separate classifiers for all choices of $k$. This observation refutes the argument that class embeddings from multiple consecutive layers that are individually strong at predicting specific classes might, when combined, introduce noise that impairs performance, which is clearly not the case in Fig. 1(b).

### v2

@@ Abstract
We address the Continual Learning (CL) problem, wherein a model must learn a sequence of tasks from non-stationary distributions while preserving prior knowledge upon encountering new experiences. With the advancement of foundation models, CL research has pivoted from the initial learning-from-scratch paradigm towards utilizing generic features from large-scale pre-training. However, existing approaches to CL with pre-trained models primarily focus on separating class-specific features from the final representation layer and neglect the potential of intermediate representations to capture low- and mid-level features, which are more invariant to domain shifts. In this work, we propose LayUP, a new prototype-based approach to continual learning that leverages second-order feature statistics from multiple intermediate layers of a pre-trained network. Our method is conceptually simple, does not require access to prior data, and works out of the box with any foundation model. LayUP surpasses the state of the art in four of the seven class-incremental learning benchmarks, all three domain-incremental learning benchmarks and in six of the seven online continual learning benchmarks, while significantly reducing memory and computational requirements compared to existing baselines. Our results demonstrate that fully exhausting the representational capacities of pre-trained models in CL goes well beyond their final embeddings. The code will be made publicly available upon acceptance.
@@ 1 Introduction
{subcaptionblock}
[t]3.2in Refer to caption {subcaptionblock}[t]3.2in Refer to caption
Figure 1: Prior works
Figure 2: LayUP
@@
Figure 3: Overview of LayUP. Prior works on CL with pre-trained models use only final representation layer features for classification. LayUP enhances the last layer representations of pre-trained models by incorporating features from intermediate layers, resulting in more accurately calibrated similarity measures for prototype-based classification and greater robustness to domain gaps in the CL setting. Stars ($\star$) denote class prototypes.
@@
Continual Learning (CL) is a subfield of machine learning dedicated to developing models capable of learning from a stream of data while striking the balance between adapting to new concepts and retaining previously acquired knowledge without forgetting (Parisi et al., 2019; Chen & Liu, 2018). While traditional works on CL primarily focus on the learning-from-scratch paradigm, the introduction of large foundation models has initiated a growing interest in developing CL methods upon the powerful representations resulting from large-scale pre-training. Continual learning with pre-trained models builds on the assumption that a large all-purpose feature extractor provides strong knowledge transfer capabilities and great robustness to catastrophic forgetting (McCloskey & Cohen, 1989) during incremental adaptation to downstream tasks.
@@
Such methods for downstream adaptation can be categorized into two general strategies (Wang et al., 2023a; McDonnell et al., 2023): (i) carefully fine-tuning the pre-trained representations (full body adaptation), or (ii) keeping the pre-trained backbone fixed while learning a small set of additional parameters such as prompts (Lester et al., 2021; Liu et al., 2023; Jia et al., 2022), adapters (Rebuffi et al., 2017a; Houlsby et al., 2019; Hu et al., 2022; Chen et al., 2022; 2023), or other methods (Ke et al., 2021; Marullo et al., 2023; Lian et al., 2022). This strategy is commonly referred to as parameter-efficient transfer learning (PETL). Both strategies (i) and (ii) have in common that they do not rely on rehearsal from past data to perform well, unlike many of the top-performing CL strategies used for models trained from scratch. Despite their merits, most approaches to CL with pre-trained models that fall under one of those strategies update a fully connected classification head (linear probe) via iterative gradient methods during training and inference, which is subject to task-recency bias (Mai et al., 2021; Rymarczyk et al., 2023; Wang et al., 2023b) and catastrophic forgetting (Zhang et al., 2023; Ramasesh et al., 2021). Class-prototype methods (Zhou et al., 2023b; Janson et al., 2022; McDonnell et al., 2023) have been established as a viable alternative, as they utilize extracted features directly, which is not only computationally efficient but also superior to CL methods applied to the trainable linear probe.
@@
Despite their effectiveness, contemporary class-prototype methods for CL extract only the features obtained from the last layer of the backbone for prototype construction. However, as the discrepancy between the pre-training and fine-tuning domains widens, the high-level features from the last layer may not generalize sufficiently to ensure adequate class separability in the target domain. Thus, the challenge lies in developing a prototype-based classifier that utilizes the representations of a pre-trained feature extractor in a holistic manner, ideally with low memory and computational costs.
@@
In this work, we hypothesize that representations are most effectively leveraged from multiple layers to construct a classifier based on first-order (class prototypes) and second-order (Gram matrix) feature statistics. Such a strategy is inspired by works on Neural Style Transfer (Gatys et al., 2016; Jing et al., 2020), which involve applying the artistic style of one image to the content of another image, a problem that can be viewed through the lens of domain adaptation (Li et al., 2017). To achieve neural style transfer, a model has to disentangle image information into content and style (e.g., textures, patterns, colors). While content information is expressed in terms of activations of features at different layers (which translates to calculating intra-layer class prototypes in our case), style information is expressed as correlations between features of different layers (which corresponds to calculating intra-layer Gram matrices).
@@
We thus propose Intra-Layer Universal Prototypes (LayUP), a class-prototype method for CL that is based on the aforementioned strategy to disentangle style and content information of images at multiple layers of a pre-trained network. An overview of our method is given in Fig. 3. LayUP obtains rich and fine-grained information about images to perform a more informed classification that is less sensitive to domain shifts. We additionally experiment with different parameter-efficient fine-tuning strategies to further refine the intra-layer representations obtained by LayUP. Across various CL benchmarks and learning settings, our method enhances performance and narrows the gap to the upper bound by as much as 80% (Stanford Cars-196 in the CIL setting) compared to the next best baseline, while also reducing its memory and compute requirements by 81% and up to 90%, respectively.
@@
Our contributions are threefold: (1) We present and examine in detail the CL strategy with pre-trained models and show why it benefits from using intra-layer representations for classification. We further demonstrate the advantages of leveraging cross-correlations between intra-layer and inter-layer features to decorrelate class prototypes. (2) Building on our insights, we propose LayUP, a novel class-prototype approach to CL that leverages second-order statistics of features from multiple layers of a pre-trained model. Inspired by prior works (Zhou et al., 2023a; McDonnell et al., 2023), we experiment with different methods for parameter-efficient model tuning and extend our method towards first session adaptation. Our final approach is not only conceptually simple, but also benefits from low memory and computational requirements, and integrates seamlessly with any pre-trained model. (3) We report performance improvements with a pre-trained Vit-B/16 (Dosovitskiy et al., 2021) backbone on the majority of benchmarks the Class-Incremental Learning (CIL) and Domain-Incremental Learning (DIL) settings (van de Ven & Tolias, 2019) as well as in the challenging Online Continual Learning (OCL) setting. We show that LayUP is especially effective under large distributional shifts and in the low-data regime. Our results highlight the importance of taking a deeper look at the intermediate layers of pre-trained models to better leverage their powerful representations, thus identifying promising directions for CL in the era of large foundation models.
@@ 2.1 Continual learning
Traditional approaches to CL consider training a model from scratch on a sequence of tasks, while striking a balance between adapting to the currently seen task and maintaining high performance on previous tasks. They can be broadly categorized into the paradigms of regularization (Kirkpatrick et al., 2017; Zenke et al., 2017; Li & Hoiem, 2018; Aljundi et al., 2018; Dhar et al., 2019), which selectively restricts parameter updates, (pseudo-)rehearsal (Rebuffi et al., 2017b; Chaudhry et al., 2019; Buzzega et al., 2020; Prabhu et al., 2020), which recovers data either from a memory buffer or from a generative model, and dynamic architectures (Rusu et al., 2016; Yoon et al., 2018; Serra et al., 2018), which allocate disjoint parameter spaces for each task.
@@ 2.2 Continual learning with pre-trained models
Leveraging the powerful representations from large foundation models for downstream continual learning has shown to not only facilitate knowledge transfer, but also to increase robustness against forgetting (Ramasesh et al., 2022; Ostapenko et al., 2022). L2P (Wang et al., 2022c) and DualPrompt (Wang et al., 2022b) outperform traditional CL baselines by training a pool of learnable parameters (i.e., prompts), which serve as queries to guide a Vision Transformer (ViT) (Dosovitskiy et al., 2021) towards adapting to downstream tasks. These approaches provide the foundation for numerous more recent prompt learning methods, such as S-Prompt (Wang et al., 2022a), DAP (Jung et al., 2023), and CODA-Prompt (Smith et al., 2023a), which further enhance performance.
Contrary to prompt learning methods that keep the backbone parameters fixed while updating a small set of additional parameters, some parallel works explore methods for making careful adjustments to the backbone parameters. SLCA (Zhang et al., 2023) uses a small learning rate for updates to the ViT backbone while training a linear probing layer with a larger learning rate. L2 (Smith et al., 2023b) applies regularization to the self-attention parameters of a ViT during continual fine-tuning. First Session Adaptation (FSA) (Panos et al., 2023) makes adjustments to the parameters of the backbone only during first task training to bridge the gap between pre-training and downstream CL domains.
As an alternative to the methods above that adjust the feature space via cross-entropy loss and backpropagation, constructing prototypes directly from the pre-trained features can avoid catastrophic forgetting during CL without storing data. Janson et al. (2022) and Zhou et al. (2023b) find that a conceptually simple Nearest Mean Classifier (NMC) outperforms various prompt learning methods. ADAM (Zhou et al., 2023b) combines NMC with FSA and concatenates the features from the initial pre-trained ViT and the adapted parameters. Other methods leverage second-order feature statistics, i.e., covariance information, for class prototype accumulation: FSA+LDA (Panos et al., 2023) applies an incremental version of linear discriminant analysis to a pre-trained ResNet encoder (He et al., 2016). RanPAC (McDonnell et al., 2023) uses high-dimensional random feature projections to decorrelate class prototypes of a pre-trained ViT backbone.
@@ 3 Preliminaries
3.1 Continual learning problem formulation
We consider a feature extractor $\phi(\cdot)$ composed of $L$ consecutive layers and a sequence of training datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$, where the $t$th task $\mathcal{D}_{t}=\{({\bm{x}}_{t,n},y_{t,n})\}_{n=1}^{N_{t}}$ contains pairs of input samples ${\bm{x}}_{t,n}\in\mathcal{X}_{t}$ and their ground-truth labels $y_{t,n}\in\mathcal{Y}_{t}$. The total number of classes seen in $\mathcal{D}$ (i.e., the number of unique elements in $\mathcal{Y}$) is denoted with $C$. We represent the $d_{l}$-dimensional encoded features from the $l$th layer of model $\phi$ from some test image ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$ with unknown label during inference as $\phi_{l}({\bm{x}})\in{\mathbb{R}}^{d_{l}}$. We consider rehearsal-free CL, where historical data cannot be fetched for replay. In the Class-Incremental Learning (CIL) setting, label spaces are partially disjoint between tasks. Conversely, in the Domain-Incremental Learning (DIL) setting, label spaces are kept across tasks, but samples within a label space are subject to distributional shifts. Online Continual Learning (OCL) is a more tightened definition of the aforementioned settings towards each datum in the stream of tasks being observed only once. All three CL settings prohibit the access of the model to task identifiers during testing (notably, our method does not require access to task-specific information during training either).
@@ 3.2 Class-prototype methods for CL with pre-trained models
To leverage the powerful representations from large-scale pre-training while overcoming the limitations of full body adaptation and linear probing, class-prototype methods construct representatives for each class from the last layer features of a pre-trained model. The most straightforward class-prototype method is the Nearest Mean Classifier (NMC), which aggregates per-class training sample features via averaging (denoted as $\overline{{\bm{c}}}_{y}$ for each class $y$) and, during inference, selects for each test sample’s feature vector the class with the smallest Euclidean distance (Janson et al., 2022) or highest cosine similarity (Zhou et al., 2023b) with its class prototype. Considering the cosine similarity measure and some ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, the predicted class label is obtained by:
@@
Although Eq. 1 and Eq. 2 are defined with respect to the maximum number of classes $C$ after observing all $T$ tasks, they can be applied after seeing any task $t\leq T$ or any sample $n\leq N_{t}$. When denoting the extracted features of some input datum ${\bm{x}}_{t,n}\in\mathcal{D}_{t}$ as ${\bm{f}}_{t,n}=\phi_{L}({\bm{x}}_{t,n})$ and considering ${\bm{F}}\in\mathbb{R}^{N\times d_{L}}$ as concatenated row-vector features ${\bm{f}}_{t,n}$ of all $N$ training samples, Eq. 3 is reduced to ${\bm{G}}={\bm{F}}^{T}{\bm{F}}$, which corresponds to the original definition of a Gram matrix as used in the closed-form ridge estimator (Hoerl & Kennard, 1970).
4 LayUP
Figure 4: Classification performance of intermediate layers. Comparison for two different pre-training paradigms of a ViT-B/16 (Dosovitskiy et al., 2021) backbone and split ImageNet-R and CIFAR-100 datasets. For each intermediate layer $l\in\{1,\dots,L-1\}$ (where $L=12$ denotes the final representation layer), the bars represent the percentage of classes for which a classifier, utilizing Eq. 2 at layer $l$, surpasses the accuracy of the classifier at the $L$-th layer.
4.1 Class prototyping from second-order intermediate features
@@
denote the concatenated input features of some input sample ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, extracted from the last $k$ layers of the pre-trained model $\phi(\cdot)$. Such features can be, e.g., class embeddings ([CLS]) of a transformer (Vaswani et al., 2017) encoder or flattened feature maps of a ResNet (He et al., 2016) encoder. This yields a modified estimator
@@
We first motivate our approach by showing that intra-layer representations capture expressive class statistics for prototype-based classification. For the split CIFAR-100 (Krizhevsky, 2009) and ImageNet-R (Hendrycks et al., 2021a) datasets and two ViT models, we construct one prototype-based classifier using Eq. 2 per layer $l\in\{1,\dots,L-1\}$ and measure the percentage of classes that layer $l$ predicts better than the last layer $L$ ($L=12$ for ViT-B/16). The results are provided in Fig. 4. Across both datasets, the classifiers from the last five intermediate layers outperform the representation layer classifier in up to 32% of the classes. The advantage of intra-layer classifiers is more pronounced in the ViT model pre-trained on 1 000 different ImageNet classes (IN1K) compared to the ViT model pre-trained on 21 000 different ImageNet classes (IN21K). The results indicate that representations derived from intermediate layers generally provide meaningful knowledge to more effectively distinguish classes at various hierarchical levels.
@@
Figure 5: Comparison of techniques to integrate intra-layer representations. LayUP implementations for different values of $k$ using a shared representation and Gram matrix as in Eq. 5 versus averaging over separate ridge (or, Gram) classifiers for each layer using Eq. 2. Results are reported as average accuracy scores over CL training on split ImageNet-R (left) and CIFAR-100 (right) datasets following phase B of Alg. 1.
@@
There are two intuitive ways to integrate intra-layer representations into the class-prototype classifier: Averaging over $k$ separate classifiers per Eq. 2 or concatenating representations from the last $k$ layers to obtain shared class prototypes per Eq. 5. Fig. 5 shows average accuracies for two split datasets and different values of $k$. For each $k$, transforming shared class-specific information via Gram matrix inversion, as described in Eq. 5, proves to be superior to averaging across separate classifiers. This finding indicates that the correlations across layers, captured in the shared Gram matrix (cf. Eq. 6), add meaningful information to enhance class separability.
While performance generally benefits from higher values of $k$, the difference in performance between $k=4$ and $k=6$ is more pronounced than that between $k=6$ and $k=12$. Considering that a higher $k$ corresponds to increased computational and memory demands as $\mathbf{G}$ and $\mathbf{c}_{y}$ increase in dimension, this substantiates $k=6$ as a reasonable choice for maximum representation depth. Although concatenating representations from multiple consecutive layers results in higher memory demands compared with using a last layer classifier or averaging over layer-wise classifiers, we will demonstrate in Sec. 4.3 that our method remains significantly lighter in memory and computational requirements than other competitive class-prototype methods for CL.
@@
Interestingly, the performance of the last-layer classifier $k=1$, as illustrated in Fig. 5, is approximately equivalent to that achieved by averaging over separate classifiers for $k=4$ and $k=6$. This observation challenges the argument that class embeddings from multiple consecutive layers, which are individually strong at predicting specific classes (cf. Fig. 4), might introduce noise that impairs performance when combined via averaging.